OpenStudy (anonymous):

A circle is described by the equation (x−12)2+(y−32)2=49. What are the cordinates for the center of the circle and the length of the radius? (-12,-32), 23units (-12,-32), 49units (12,32), 49units (12,32), 23units

OpenStudy (nincompoop):

do you know the general form for circle?

OpenStudy (anonymous):

\[\large (x-h)^{2}~+~(y-k)^{2}=r^{2}\] \[\large (h,k)~=~Center~coordinates~of~circle\] \[\large r~=~radius\]

OpenStudy (nincompoop):

that's a lot of letters and where do they exist in a graph?

OpenStudy (anonymous):

So if we break down each part it will look like this \[\large (x-h)^{2}~=~(x-12)^{2}\] From that, what would h be? \[\large (y-k)^{2}~=~(y-32)^{2}\] From that, what would k be? \[\large r^{2}~=~49\] From that, what would r be?